# How Similar Are Grokipedia and Wikipedia? A Multi-Dimensional Textual and Structural Comparison

### v2

@@ Abstract
The launch of Grokipedia, an AI-generated encyclopedia developed by Elon Musk’s xAI, was presented as a response to perceived ideological and structural biases in Wikipedia, aiming to produce “truthful” entries via the large language model Grok. Yet whether an AI-driven alternative can escape the biases and limitations of human-edited platforms remains unclear. This study undertakes a large-scale computational comparison of 1,800 matched article pairs between Grokipedia and Wikipedia, drawn from the 2,000 most-edited Wikipedia pages. Using metrics across lexical richness, readability, structural organization, reference density, and semantic similarity, we assess how closely the two platforms align in form and substance. The results show that while Grokipedia exhibits strong semantic and stylistic alignment with Wikipedia, it typically produces longer but less lexically diverse articles, with fewer references per word and greater structural variability. These findings suggest that AI-generated encyclopedic content currently mirrors Wikipedia’s informational scope but diverges in editorial norms, favoring narrative expansion over citation-based verification. The implications highlight new tensions around transparency, provenance, and the governance of knowledge in an era of automated text generation.
@@ 1 Introduction
Online encyclopedias have become infrastructural to public knowledge, shaping how people learn about politics, science, culture, and current events. Wikipedia, launched in 2001, remains the paradigmatic example: a nonprofit, volunteer-edited project governed by policies such as Neutral Point of View (NPOV) and verifiability. Despite its success, Wikipedia has long faced questions about bias, reliability, and systemic underrepresentation. [^1] For a survey of evidence and debates on reliability and bias, see e.g., [4, 8, 17, 18].
@@
In late 2025, xAI introduced Grokipedia, an AI-generated encyclopedia positioned as an explicit alternative to Wikipedia. [^2] See [20, 1, 19, 11, 13, 6, 10]. According to xAI’s framing, Grokipedia aims to “purge out the propaganda” and provide “truthful” entries, with content generated and internally “fact-checked” by the Grok language model rather than curated by a community of human editors.[20, 11, 10] At launch (October 27, 2025), Grokipedia reportedly contained roughly 800–900 thousand entries, supported a suggest-edit rather than direct-edit workflow, and offered only limited transparency about licensing and code provenance.[20, 19, 1] Early commentary focused on two recurring issues: (i) claims that many Grokipedia pages were copied or closely adapted from Wikipedia, and (ii) concerns about ideological bias, hallucinated citations, and other generative artifacts typical of large language models.[13, 6, 11, 10]
@@
The launch has reignited a longstanding question: is Wikipedia biased, and if so, can an AI-first encyclopedia do better? Empirically, prior research documents multiple bias forms in Wikipedia, early left–right slant in political pages that attenuated over time,[4] topical and coverage imbalances,[8] and framing asymmetries across ideology and gender.[17, 18] The dynamics of editorial conflict and consensus formation have also been studied extensively, showing that edit activity, controversy, and attention are strongly correlated.[26] Parallel work on large language models (LLMs) shows measurable political and cultural biases that vary across model families and prompts,[22, 15, 3, 9] raising distinct questions about the epistemic reliability of AI-generated text.
@@
Against this backdrop, our study asks a concrete, testable question: for topics that exist on both platforms, how similar are Grokipedia and Wikipedia in practice? We address this by pairing articles across platforms and comparing them along four analytic dimensions: (1) lexical similarity (TF–IDF cosine, unigram and $n$-gram overlap), (2) semantic similarity (embedding cosine and BERTScore), (3) structural similarity (section and paragraph organization), and (4) stylistic similarity (readability, lexical diversity, part-of-speech composition). In addition, we examine descriptive indicators, article length, readability, reference density, and link structure, since editorial workflows (human versus model-generated) shape article form as much as content.
@@
To construct a broad and representative corpus, we began with Wikipedia’s 2,000 most-edited English-language articles, a population empirically linked to controversy and social salience.[26] Among these, approximately 1,800 titles had valid, nontrivial matches on Grokipedia. This enlarged sample enables more stable estimates of cross-platform differences than earlier pilot comparisons limited to a few hundred pairs.
@@
This comparison provides empirical grounding for evaluating the distinctiveness, or sameness, of Grokipedia relative to Wikipedia. If Grokipedia largely reproduces Wikipedia’s textual and structural patterns, the platform may represent an AI-mediated continuity rather than a substantive editorial departure. Conversely, systematic differences in length, style, or citation density could reveal an emerging form of automated editorial logic, distinct from Wikipedia’s norms of transparency, consensus, and verifiability. By quantitatively characterizing these dimensions across matched topics, we aim to clarify where the two encyclopedias align, where they diverge, and what those patterns imply for neutrality, provenance, and the evolving governance of reference knowledge in the age of generative AI. [^3] For broader debates on neutrality, transparency, and provenance in AI-generated reference content, see [11, 6, 10, 23].
@@ 2.1 Sampling and Data Collection
We analyzed the 2,000 most-edited English-language Wikipedia articles as of October 2025, identified via cumulative edit counts. Prior research shows that heavily edited entries correlate strongly with controversy, topical salience, and social polarization [26]. This sampling strategy thus prioritizes articles that are both textually substantial and socially significant.
For each title, we retrieved the corresponding entries from Wikipedia and Grokipedia, generating URLs of the form https://en.wikipedia.org/wiki/<Title>
and
https://Grokipedia.com/page/<Title>. HTML pages were downloaded via the requests library (Python 3.11) with polite delays and a standard user-agent header, then parsed using BeautifulSoup4. The extraction and analysis pipeline continuously wrote progress files (wiki_Grokipedia_similarity_live.csv) to enable incremental recovery and monitoring. After host-aware text extraction (see below), only pairs in which both pages contained at least 500 words of clean prose were retained. Of the 2,000 target titles, 1,811 matched article pairs satisfied these criteria and formed the final analytical sample.
@@ 2.2 Text Cleaning and Feature Extraction
We implemented a host-aware parsing strategy tailored to each platform’s HTML structure to maximize content fidelity.
@@
Each cleaned article was tokenized into sentences and words using nltk’s Punkt tokenizer. A comprehensive set of descriptive, structural, and stylistic metrics was computed:
  • Structural features: counts of paragraphs, headings (h1–h4), hyperlinks, images, and references; visible character counts; and derived ratios such as references, links, and headings per 1,000 words.
@@ 2.3 Similarity Measures Between Platforms
To quantify cross-platform alignment, we computed a suite of nine similarity measures grouped into four conceptual domains:
@@
    Semantic similarity: cosine similarity between SentenceTransformer embeddings (all-MiniLM-L6-v2[14], and contextual similarity via BERTScore F1 for the first 50 sentences per article [27].
@@
    Structural and stylistic similarity: a composite of (a) overlap in detected section headings and (b) proportional similarity of median paragraph lengths (weighted 0.6 and 0.4, respectively), together with normalized Manhattan distance across stylistic profiles (sentence length, lexical diversity, readability, POS composition) transformed to a 0–1 scale.
A composite similarity index was calculated as the mean of lexical, semantic, structural, and stylistic sub-scores, providing a general indicator of cross-platform convergence.
@@ 2.4 Statistical and Comparative Analysis
All analyses were performed in Python 3.11 using pandas, numpy, and scikit-learn. Descriptive statistics (mean $\pm$ SD) were calculated for all features. Differences between Grokipedia and Wikipedia were evaluated using paired $t$-tests for normally distributed measures and Wilcoxon signed-rank tests otherwise. Spearman rank correlations among similarity metrics were computed to assess interdependence and clustering. Visualizations, including histograms, correlation heatmaps, and mean $\pm$ SE summary plots, were generated with matplotlib and exported directly to PDF for inclusion in the manuscript.
@@ 2.5 Limitations
The dataset covers only topics that are both heavily edited on Wikipedia and available on Grokipedia. Because edit frequency correlates with controversy and visibility [26], the sample likely overrepresents politically or culturally salient topics. In addition, Grokipedia’s content-generation pipeline is opaque, potentially introducing stylistic inflation (e.g., verbosity, hallucinated references) or structural artifacts that differ from human-edited prose. Nevertheless, the scale of the dataset and the matched-pair design provide a robust basis for assessing how human- and AI-generated encyclopedic text diverge in length, structure, and informational alignment.
@@ 3.1 Sample and Coverage
From the expanded list of the 2,000 most-edited English Wikipedia titles, we successfully retrieved and matched 1,811 Grokipedia–Wikipedia article pairs that passed quality checks (both pages available and $\geq$500 cleaned words each). [^4] Clean word counts were computed from visible prose only, excluding infoboxes, scripts, and navigation templates. This larger corpus substantially extends the coverage of the original 416-title sample while preserving topical breadth and diversity.
@@ 3.2 Descriptive Differences Between Platforms
Figure 1 compares article-length distributions on a logarithmic scale. Grokipedia entries remain systematically longer than their Wikipedia counterparts, though the gap narrows compared with the smaller sample. Readability, measured by the Flesch–Kincaid grade level (panel B of Figure 2), is again higher on Grokipedia, indicating more syntactically complex and less reader-friendly prose.
Table 1 summarises key descriptive statistics. Across the enlarged dataset, Grokipedia articles average about 14,200 words versus 9,400 on Wikipedia, and they exhibit lower lexical diversity and far fewer explicit references, links, and headings per thousand words. These contrasts reinforce that Grokipedia expands text length through elaboration rather than citation density or structural detail. Figure 2 visualises these platform differences (means $\pm$ SE) across all six measures reported in Table 1.
@@
Table 1: Descriptive statistics across matched article pairs (N=1,811N{=}1{,}811). Values are means [SD]; $p$ from paired $t$-tests.
@@
Clean word count 14,200 [5,650] 9,360 [4,710] +4,840 $10^{-200}$
Flesch–Kincaid grade 19.3 [6.6] 12.8 [9.9] +6.5 $10^{-100}$
Lexical diversity (TTR) 0.244 [0.057] 0.248 [0.049] -0.004 0.02
References per 1k words 21.4 [7.4] 90.2 [218] -68.8 $10^{-50}$
Links per 1k words 16.8 [6.0] 428 [683] -411 $10^{-120}$
Headings (H2–H4)/1k words 2.87 [1.20] 5.58 [8.32] -2.71 $10^{-40}$
@@ 3.3 Cross-Platform Similarity
To assess how closely Grokipedia mirrors Wikipedia, we computed lexical, semantic, structural, and stylistic similarity metrics for each article pair. Figure 3 shows the correlation structure among these metrics: lexical and semantic measures form a tight cluster, while structural and stylistic dimensions vary more independently. Table 2 reports the average values across all pairs.
Figure 3: Correlation among similarity metrics (Pearson’s $r$) across 1,811 article pairs.
Table 2: Average similarity metrics across matched pairs (N=1,811N{=}1{,}811).
Metric Mean [SD]
TF–IDF cosine 0.505 [0.149]
Jaccard (unigram) 0.362 [0.217]
Overlap (1-gram) 0.597 [0.213]
Overlap (2-gram) 0.329 [0.339]
Overlap (3-gram) 0.250 [0.374]
Semantic cosine 0.825 [0.106]
BERTScore (F1) 0.193 [0.334]
Structural similarity 0.216 [0.163]
Stylistic similarity 0.835 [0.072]

### v3

@@ Abstract
The launch of Grokipedia—an AI-generated encyclopedia developed by Elon Musk’s xAI—was presented as a response to perceived ideological and structural biases in Wikipedia, aiming to produce “truthful” entries using the Grok large language model. Yet whether an AI-driven alternative can escape the biases and limitations of human-edited platforms remains unclear. This study conducts a large-scale computational comparison of more than 17,000 matched article pairs from the 20,000 most-edited English Wikipedia pages. Using metrics spanning lexical richness, readability, reference density, structural features, and semantic similarity, we assess how closely the two platforms align in form and substance. We find that Grokipedia articles are substantially longer and contain significantly fewer references per word. Moreover, Grokipedia’s content divides into two distinct groups: one that remains semantically and stylistically aligned with Wikipedia, and another that diverges sharply. Among the dissimilar articles, we observe a systematic rightward shift in the political bias of cited news sources, concentrated primarily in entries related to politics, history, and religion. These findings suggest that AI-generated encyclopedic content diverges from established editorial norms—favouring narrative expansion over citation-based verification. The implications highlight emerging tensions around transparency, provenance, and the governance of knowledge in an era of automated text generation.
@@ 1 Introduction
Online encyclopedias have become infrastructural to public knowledge, shaping how people learn about politics, science, culture, and current events. Wikipedia, launched in 2001, remains the paradigmatic example: a nonprofit, volunteer-edited project governed by policies such as Neutral Point of View (NPOV) and verifiability. Despite its success, Wikipedia has long faced questions about bias, reliability, and systemic underrepresentation. [^1] For a survey of evidence and debates on reliability and bias, see e.g., [GreensteinZhu2012AER, ReliabilityWikipedia2019, MIWikipediaBias2024, RightingWriters2025].
@@
In late 2025, xAI introduced Grokipedia, an AI-generated encyclopedia positioned as an explicit alternative to Wikipedia. [^2] See [GrokipediaWP, AP2025, BI2025, Guardian2025, Verge2025, Wired2025, Time2025]. According to xAI’s framing, Grokipedia aims to “purge out the propaganda” and provide “truthful” entries, with content generated and internally “fact-checked” by the Grok language model rather than curated by a community of human editors.[GrokipediaWP, Guardian2025, Time2025] At launch (October 27, 2025), Grokipedia reportedly contained roughly 800-900 thousand entries, supported a suggest-edit rather than direct-edit workflow, and offered only limited transparency about licensing and code provenance.[GrokipediaWP, BI2025, AP2025] Early commentary focused on two recurring issues: (i) claims that many Grokipedia pages were copied or closely adapted from Wikipedia, and (ii) concerns about ideological bias, hallucinated citations, and other generative artifacts typical of large language models.[Verge2025, Wired2025, Guardian2025, Time2025]
@@
The launch has reignited a longstanding question: is Wikipedia biased, and if so, can an AI-first encyclopedia do better? Empirically, prior research documents multiple bias forms in Wikipedia, early left–right slant in political pages that attenuated over time,[GreensteinZhu2012AER] topical and coverage imbalances,[ReliabilityWikipedia2019] and framing asymmetries across ideology and gender.[MIWikipediaBias2024, RightingWriters2025] The dynamics of editorial conflict and consensus formation have also been studied extensively, showing that edit activity, controversy, and attention are strongly correlated.[yasseri2012dynamics]
Parallel research on large language models (LLMs) demonstrates measurable political and cultural biases that vary across model architectures and prompts [YangMenczer2023, Rettenberger2024, Exler2025, Jenny2023]. LLMs exhibit human-like biases based on input framing [lior2025wildframe] and show intrinsic political biases that affect text continuation and prediction [lin2024investigating], raising questions about the epistemic reliability of AI-generated content. Grokipedia, in particular, relies on the Grok LLM developed by xAI. In comparative studies on bibliographic reference retrieval, Grok stood out as one of the few models that did not fabricate references, although roughly 40% of references contained incomplete or inaccurate details, such as publication year, journal name, or author information [sidorenko2025assessing]. Similarly, in another study, researchers evaluated four LLMs, including Grok 2, to examine how source framing influences model judgments. Their results indicate that Grok 2 demonstrates high consistency and minimal ideological bias, with only modest shifts in evaluation when the framing of the input changed [germani2025source]. These results suggest that Grok may have the potential to generate reliable and neutral encyclopedic content, although intrinsic biases in LLMs could still persist, raising the question of whether Grok can produce a less biased encyclopedia than Wikipedia.
@@
Against this backdrop, our study asks a concrete, testable question: for topics that exist on both platforms, how similar are Grokipedia and Wikipedia in practice? We address this by pairing articles across platforms and comparing them along three similarity dimensions: (1) lexical similarity (TF–IDF cosine, unigram and $n$-gram overlap), (2) semantic similarity (embedding cosine and BERTScore), and (3) stylistic similarity (readability, lexical diversity, part-of-speech composition). In addition, we examine descriptive indicators, article length, readability, reference density, and link structure, since editorial workflows (human versus model-generated) shape article form as much as content.
We also measure the political bias of each article by aggregating the bias scores of its cited references (see Methods for details). To contextualize these patterns, we classify all articles into topical domains and then analyse how both similarity and the shift in political bias vary across topics.
@@
To construct a broad and representative corpus, we began with Wikipedia’s 20,000 most-edited English-language articles, a population empirically linked to controversy and social salience.[yasseri2012dynamics] Among these, approximately 17,700 titles had valid, nontrivial matches on Grokipedia.
@@
This comparison provides empirical grounding for evaluating the distinctiveness, or sameness, of Grokipedia relative to Wikipedia. If Grokipedia largely reproduces Wikipedia’s textual and structural patterns, the platform may represent an AI-mediated continuity rather than a substantive editorial departure. Conversely, systematic differences in length, style, or citation density could reveal an emerging form of automated editorial logic, distinct from Wikipedia’s norms of transparency, consensus, and verifiability. By quantitatively characterizing these dimensions across matched topics, we aim to clarify where the two encyclopedias align, where they diverge, and what those patterns imply for neutrality, provenance, and the evolving governance of reference knowledge in the age of generative AI. [^3] For broader debates on neutrality, transparency, and provenance in AI-generated reference content, see [Guardian2025, Wired2025, Time2025, GrogConversation2025].
@@ 2.1 Sampling and Data Collection
We analyzed the 20,000 most-edited English-language Wikipedia articles as of October 2025, identified via cumulative edit counts. Prior research shows that heavily edited entries correlate strongly with controversy, topical salience, and social polarization [yasseri2012dynamics]. To ensure that our comparison focused on substantive articles, we excluded all list-style pages as well as titles that were date- or year-like rather than topical. Specifically, we removed (a) pages with the title ”List of Topic” (e.g., UFC champions) (b) calendar dates (e.g., March 11), (c) standalone years (1000–2099), (d) “Year in Topic” pages (e.g., 2003 in music), (e) “Deaths in Month Year” pages, and (f) standalone month names.
For each remaining title, we retrieved the corresponding entries from Wikipedia and Grokipedia, generating URLs of the form https://en.wikipedia.org/wiki/<Title> and
https://Grokipedia.com/page/<Title>.
HTML pages were downloaded between 5–11 November 2025 using the requests library (Python 3.11) with polite delays and a standard user-agent header, and then parsed using BeautifulSoup4. After host-aware text extraction (described below), we retained only article pairs in which both platforms produced at least 500 words of clean prose. Of the original 20,000 target titles, 17,790 matched pairs met these criteria and formed the final analytical sample.
@@ 2.2 Text Cleaning and Feature Extraction
We implemented a host-aware parsing strategy tailored to each platform’s HTML structure to maximise content fidelity.
@@
Each cleaned article was tokenized into sentences and words using nltk’s Punkt tokenizer. A comprehensive set of descriptive, and stylistic metrics was computed.
@@ 2.3 Similarity Measures Between Platforms
To quantify cross-platform alignment, we computed a suite of eight similarity measures grouped into four conceptual domains:
@@
    Semantic similarity: cosine similarity between SentenceTransformer embeddings (all-MiniLM-L6-v2[reimers2019sentence], and contextual similarity via BERTScore F1 for the first 50 sentences per article [zhang2019bertscore].
@@
    stylistic similarity: The Manhattan distance across stylistic profiles (sentence length, lexical diversity, readability, POS composition) transformed to a 0–1 scale.
To verify the extent to which these metrics reflected a single latent construct, we conducted a principal components analysis (PCA) on the full set of similarity measures. The first principal component accounted for the vast majority of total variance, with all measures loading in the expected (positive) direction. This indicated that the metrics were largely co-linear and captured a common underlying dimension of article similarity. Based on this result, we constructed a unified combined similarity score, defined as the standardized value of the first principal component, which we use in all subsequent analyses.
2.4 Political Bias ratings
Outlet-level political-leaning scores were taken from the News Media Bias and Factuality dataset [sanchez2024mapping], mapped to a symmetric numeric scale, and used only when a numeric bias was available for a domain or its brand variant (e.g., bbc.com/bbc.co.uk). Non-news and reference/utility hosts (e.g., doi.org, archive.org, worldcat.org, major platforms, and *.gov/*.edu) were excluded. For brand merging we used eTLD+1 normalization plus a brand key (leftmost label).
For each article, bias shift was computed as the mention-weighted mean outlet bias in Grokipedia minus the corresponding mean in Wikipedia (positive values indicate a rightward shift; negative values indicate a leftward shift). Centers of mass in Figure 6 were derived using Gaussian kernel density estimation (KDE) applied to the raw article-level points, computed separately for the low- and high-similarity partitions. For the pie charts in Figure 5, we restricted the analysis to the low-similarity subset (articles with a negative combined similarity score) and plotted the top 40% cumulative share of referenced news outlets. Slice colors encode outlet bias, and panel titles report the mention-weighted average platform bias normalized to ±100%.
2.5 Topics
To categorize the topics of the articles, we developed a multi-stage large-language-model (LLM) classification pipeline. We first constructed a fixed taxonomy of fourteen semantically distinct categories (e.g., History, Politics and Political Figures, Science and Technology, Literature and Art). Article titles were processed in batches of up to 300 and passed to the GPT-5-nano-2025-08-07 model with strict instructions to assign each title to exactly one category and to avoid the fallback use of “Other” whenever a reasonable category was available. The model returned a JSON-formatted list of titles grouped by category, and the script automatically retried up to five times if the output contained formatting errors. All valid batch outputs were then combined into an initial classification table mapping each title to its assigned category.
Because LLM-generated labels can occasionally be inconsistent across runs, we used a multi-round refinement procedure to improve reliability. The entire classification process was repeated across several independent rounds using the same model and prompt. After each round, titles labeled “Other” were separated and given to the model for reclassification. We then merged the rounds. Any article still lacking a category after all rounds was assigned “Other”. This multi-round integration approach increased coverage, reduced noise, and yielded a more stable and robust set of topic labels for all article pairs.
Articles were assigned to one of the following categories:
  • History
  • Politics and Political Figures
  • Geography, Nations, and Cities
  • Music, Movies, and Entertainment Celebrities
  • Sports, Athletes, and Teams
  • Science, Technology, and Knowledge
  • Religion and Ideologies
  • Brands and Products
  • Languages
  • Literature and Art
  • Health and Environment
  • Business and Infrastructure
  • Animals and Nature
  • Other
@@ 2.6 Statistical and Comparative Analysis
All analyses were performed in Python 3.11 using pandas, numpy, and scikit-learn. Descriptive statistics (mean $\pm$ SD) were calculated for all features. Differences between Grokipedia and Wikipedia were evaluated using paired $t$-tests for normally distributed measures and Wilcoxon signed-rank tests otherwise. Spearman rank correlations among similarity metrics were computed to assess interdependence and clustering. Visualizations, including histograms, correlation heatmaps, and mean $\pm$ SE summary plots, were generated with matplotlib.
@@ 2.7 Limitations
The dataset covers only topics that are both heavily edited on Wikipedia and available on Grokipedia. Because edit frequency correlates with controversy and visibility [yasseri2012dynamics], the sample likely overrepresents politically or culturally salient topics. In addition, Grokipedia’s content-generation pipeline is opaque, potentially introducing stylistic inflation (e.g., verbosity, hallucinated references) or structural artifacts that differ from human-edited prose. Nevertheless, the scale of the dataset and the matched-pair design provide a robust basis for assessing how human- and AI-generated encyclopedic text diverge in length, structure, and informational alignment.
@@ 3.1 Sample and Coverage
From the list of the 20,000 most-edited English Wikipedia articles, we successfully retrieved and matched 17,790 Grokipedia–Wikipedia article pairs that passed all quality checks (see more details in Methods). This corpus provides substantial coverage across a wide and diverse range of topics.
@@ 3.2 Descriptive Differences Between Platforms
Figure 1 compares the distribution of article lengths across platforms on a logarithmic scale. Overall, Grokipedia entries are systematically longer than their Wikipedia counterparts. While Wikipedia contains a larger number of very short articles, Grokipedia articles exhibit a pronounced peak around 7,000 words, indicating that most Grokipedia entries are substantially more verbose. Readability, measured by the Flesch–Kincaid grade level (Figure 2 (b)) and lexical diversity (Figure 2 (c), is higher on Grokipedia, indicating more syntactically complex and less reader-friendly prose.
Table 1 summarises key descriptive statistics. Across the dataset, Grokipedia articles average 7,662 words versus 6,280 on Wikipedia, and they exhibit far fewer explicit references, links, and headings per thousand words. These contrasts reinforce that Grokipedia expands text length through elaboration rather than citation density or structural detail. Figure 2 visualises these platform differences (means $\pm$ SE) across all six measures reported in Table 1.
@@
Table 1: Descriptive differences between Grokipedia and Wikipedia articles. Means and standard deviations (in brackets) are reported for each measure. $\Delta$ denotes the mean difference (Grokipedia - Wikipedia), and $p$ reports two-sided paired comparison $p$-values
@@
Clean word count 7662.36 [3662.22] 6280.51 [3828.40] 1381.846 $10^{-10}$
Flesch–Kincaid grade 14.53 [3.93] 10.67 [1.69] 3.859 $10^{-10}$
Lexical diversity (TTR) 0.29 [0.05] 0.28 [0.06] 0.006 $10^{-10}$
References per 1k words 20.78 [9.86] 35.03 [37.06] -14.257 $10^{-10}$
Links per 1k words 24.18 [16.14] 356.58 [323.37] -332.405 $10^{-10}$
Headings (H2–H4)/1k words 4.52 [1.97] 6.17 [5.19] -1.653 $10^{-10}$
@@ 3.3 Cross-Platform Similarity
To assess how closely Grokipedia mirrors Wikipedia, we computed lexical, semantic, and stylistic similarity metrics for each article pair. Figure 3 shows the correlation structure among these metrics: lexical and semantic measures form a tight cluster, while stylistic dimensions vary more independently. Table 2 reports the average values across all pairs.
Figure 3: Correlation among similarity metrics (Pearson’s $r$) across 1,8287 article pairs.
Table 2: Means and standard deviations (in brackets) are listed for each similarity metric, along with observed minimum and maximum values. Higher scores indicate greater similarity between Grokipedia and Wikipedia articles. (N=17,790N{=}17{,}790)
Metric Mean [SD] Min Max
TF–IDF cosine 0.684 [0.186] 0.024 1.000
Jaccard (unigram) 0.434 [0.299] 0.046 1.000
Overlap (1-gram) 0.640 [0.235] 0.190 1.000
Overlap (2-gram) 0.411 [0.373] 0.048 1.000
Overlap (3-gram) 0.373 [0.412] 0.005 1.000
Semantic cosine 0.862 [0.102] 0.160 1.000
BERTScore (F1) 0.306 [0.369] -0.606 0.983
Stylistic similarity 0.882 [0.076] 0.429 0.999